OpenStudy (anonymous):

18. Solve 3x – 2 = –11 for x. (Points : 1) -2 -3 2 3 19. Solve x/14 + 3 = –1 for x. (Points : 1) -55 -56 -57 -58 20. Solve x/14 – 3 = –1 for x. (Points : 1) 28 30 32 34

OpenStudy (anonymous):

18. 3x-2=-11 3x= -9 x=-9/3 x=-3 19. x/14 +3 = -1 x/14=-4 x=-4*14 x= -56 20. same as in 19th